# Wall in the stability space of the gluing stability conditions on Hirzebruch surfaces

### v1

@@ 1.7. Acknowledgement
The author would like to express his gratitude to Takayuki Uchiba for his valuable advice and numerous enlightening discussions. The author also thanks Yuki Matsubara for his extensive discussions and many helpful suggestions.
@@ 2.1. Stability conditions
    ${\rm inf}\left\{\frac{|Z(E)|}{\|v(E)\|}\ \middle|\ E\neq 0\in\mathcal{P}(\phi),\phi\in\mathbb{R}\right\}\geq C$ (2.1)
@@ 2.3. Inducing stability conditions
$\textrm{Hom}_{\mathcal{D}^{\prime}}(F(A),F(B))=0\implies\textrm{Hom}_{\mathcal{D}}(A,B)=0.$ (2.4)
@@ 2.6. Gluing construction on Ruled surfaces
$\textrm{Stab}(\mathcal{D}_{1})=\left\{(Z_{1},\mathcal{P}_{1})\ \middle|\ \begin{split}&(Z,\mathcal{P})\in\textrm{Stab}(C),Z_{1}=Z\circ F_{1}^{-1},\\ &\mathcal{P}_{2}(\phi)=p^{*}\mathcal{P}(\phi)\otimes\mathcal{O}_{S}(-C_{0})\ {\rm for\ all}\ \phi\in\mathbb{R}\end{split}\right\},$ (2.5)
$\textrm{Stab}(\mathcal{D}_{2})=\left\{(Z_{2},\mathcal{P}_{2})\ \middle|\ \begin{split}&(Z,\mathcal{P})\in\textrm{Stab}(C),Z_{2}=Z\circ F_{2}^{-1},\\ &\mathcal{P}_{2}(\phi)=p^{*}\mathcal{P}(\phi)\ {\rm for\ all}\ \phi\in\mathbb{R}.\end{split}\right\},$ (2.6)
@@
$F_{1}:K(D^{b}(C))\cong K(\mathcal{D}_{1}),\quad F_{2}:K(D^{b}(C))\cong K(\mathcal{D}_{2})$ (2.7)
@@ 3. Construction
$D^{b}(\Sigma_{e})=\langle p^{*}D^{b}(\mathbb{P}^{1})\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0}),p^{*}D^{b}(\mathbb{P}^{1})\rangle.\\$ (3.1)
@@ 3.1. Gluing construction on $D^{b}(\Sigma_{e})$
$\textrm{Hom}_{D^{b}(\Sigma_{e})}(p^{*}E_{1}\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0}),p^{*}E_{2}))\cong\textrm{Hom}_{D^{b}(\mathbb{P}^{1})}(E_{1},E_{2}\otimes(\mathcal{O}\oplus\mathcal{O}(-e))).$
@@
$E^{\bullet}:=\cdots\to 0\to\mathcal{O}(m_{j^{\prime}-1})^{\oplus n_{j^{\prime}-1}}\to\mathcal{O}(m_{j^{\prime}})^{\oplus n_{j^{\prime}}}\to 0\to\cdots$ (3.2)
@@
$\bigoplus_{i\leq 0}\textrm{Ext}^{i}(\mathcal{O}(k-1)[j_{1}+1],E^{\bullet}[j_{2}])=\bigoplus_{i\leq 0}\textrm{Ext}^{i}(\mathcal{O}(k)[j_{1}],E^{\bullet}[j_{2}])=0.$ (3.3)
@@
$\textrm{Ext}^{i+j_{2}-(j_{1}+1)}(\mathcal{O}(k-e-1),\mathcal{F})\quad and\quad\textrm{Ext}^{i+j_{2}-j_{1}}(\mathcal{O}(k-e),\mathcal{F})$ (3.4)
@@
$\mathcal{O}(k)^{\oplus i-1-n}\to\mathcal{O}(n)[1]\to\mathcal{O}(k-1)[1]^{\oplus i-n}.$ (3.5)
@@
$\mathcal{O}(k)\to\mathcal{O}_{x}\to\mathcal{O}(k-1)[1]$ (3.6)
@@
    $\sigma_{k,\zeta_{0},\zeta_{1}}:=(Z_{\zeta_{0},\zeta_{1}},\mathcal{A}(k)[j]),$ (3.8)
@@
    $Z_{\zeta_{0},\zeta_{1}}(\mathcal{O}(k-1)[j+1])=\zeta_{0},\quad Z_{\zeta_{0},\zeta_{1}}(\mathcal{O}(k)[j])=\zeta_{1}.$ (3.9)
@@
$r_{\lambda_{1}}=\textrm{ch}_{0}(\lambda_{1}(F)),\quad d_{\lambda_{1}}=\textrm{ch}_{1}(\lambda_{1}(F)),\quad r_{\rho_{2}}=\textrm{ch}_{0}(\rho_{2}(F)),\quad d_{\rho_{2}}=\textrm{ch}_{1}(\rho_{2}(F)).$ (3.10)
@@
Next, we show that these four types of pre-stability conditions satisfy the support property. To this end, we first require some preliminary results for the cases $m=1,2,3$. In the cases $m=1$ and $3$ (where $\sigma_{2}$ is a quiver stability condition), the following criterion applies.
@@
$|Z_{1}(E)|\leq C|Z_{gl,2}(E)|\quad\textrm{and}\quad|Z_{2}(E)|\leq(1+C)|Z_{gl,2}(E)|$ (3.11)
@@
We first consider the cases $m=1$ and $3$ (where $\mathcal{A}_{2}$ is the heart of a quiver stability condition). We apply the criterion in Lemma 3.10. Let $E\in\mathcal{A}_{2}$ be a non-zero object. Its central charge is given by $Z_{2}(E)=n_{0}\zeta_{0}+n_{1}\zeta_{1}$, where $\zeta_{0},\zeta_{1}\in\mathbb{H}$, and the dimension vector $(n_{0},n_{1})$ consists of non-negative integers. Since $E\neq 0$, the dimension vector is non-zero, meaning at least one of $n_{0}$ or $n_{1}$ is a positive integer. Observe that the phase of the sum $n_{0}\zeta_{0}+n_{1}\zeta_{1}$ is bounded by the phases of $\zeta_{0}$ and $\zeta_{1}$. We define $\theta:=\min\{\phi(\zeta_{0}),\phi(\zeta_{1})\}$. Since $\zeta_{0},\zeta_{1}\in\mathbb{H}$, the image $Z_{2}(\mathcal{A}_{2}\setminus\{0\})$ is contained in the sector $\mathbb{H}_{\theta}=\{re^{i\pi\phi}\mid\phi\in[\theta,1]\}$. Thus, the condition of Lemma 3.10 is satisfied, and the support property holds.
@@
Finally, for the case $m=4$, Lemma 4.2 and Proposition 4.1 imply that $\sigma$ is divisorial stability condition. Specifically, there exist divisors $B,\omega\in\textrm{NS}(S)_{\mathbb{R}}$ with $\omega$ ample such that the stability condition $\sigma_{gl,4}$ can be written as
$\sigma_{gl,4}=(Z_{\omega,B},\mathcal{A}_{\omega,B})$
(cf. (4.2)). Consequently, the support property follows from the BG-type inequality established in [MS17] Theorem 6.13.
The last statement follows from the well-known fact that any stability condition satisfying the support property is locally finite. ∎
@@
$C=\begin{pmatrix}-1&1\\ 1-k&k\end{pmatrix}$. (3.13)
@@
$C^{-1}=\begin{pmatrix}-k&1\\ 1-k&1\end{pmatrix}$. (3.14)
@@
$\begin{pmatrix}\textrm{rank}\ E\\ \textrm{deg}\ E\end{pmatrix}=C\begin{pmatrix}n_{0}\\ n_{1}\end{pmatrix}=\begin{pmatrix}n_{1}-n_{0}\\ (1-k)n_{0}+kn_{1}\end{pmatrix}$. (3.15)
@@
$\begin{pmatrix}n_{0}\\ n_{1}\end{pmatrix}=\begin{pmatrix}-k&1\\ 1-k&1\end{pmatrix}\begin{pmatrix}r_{*}\\ d_{*}\end{pmatrix}=\begin{pmatrix}d_{*}-kr_{*}\\ d_{*}+(1-k)r_{*}\end{pmatrix}$. (3.16)
@@
$+(-1)^{j_{2}}((d_{\rho_{2}}-kr_{\rho_{2}})\zeta^{\prime}_{0}+(d_{\rho_{2}}+(1-k)r_{\rho_{2}})\zeta^{\prime}_{1}).$ (3.17)

### v2

@@ 1.7. Acknowledgement
The author would like to express his gratitude to Takayuki Uchiba for his valuable advice and numerous enlightening discussions. The author also thanks Yuki Matsubara for his extensive discussions and many helpful suggestions. The author is grateful to Tomohiro Karube for his precise remarks regarding the support property. Thanks are also due to Franco Rota for reading the manuscript and providing useful comments.
@@ 2.1. Stability conditions
    (2.1) ${\rm inf}\left\{\frac{|Z(E)|}{\|v(E)\|}\ \middle|\ E\neq 0\in\mathcal{P}(\phi),\phi\in\mathbb{R}\right\}\geq C$
@@ 2.3. Inducing stability conditions
(2.5) $\textrm{Hom}_{\mathcal{D}^{\prime}}(F(A),F(B))=0\implies\textrm{Hom}_{\mathcal{D}}(A,B)=0.$
@@ 2.6. Gluing construction on Ruled surfaces
(2.6) $\textrm{Stab}(\mathcal{D}_{1})=\left\{(Z_{1},\mathcal{P}_{1})\ \middle|\ \begin{split}&(Z,\mathcal{P})\in\textrm{Stab}(C),Z_{1}=Z\circ F_{1}^{-1},\\ &\mathcal{P}_{2}(\phi)=p^{*}\mathcal{P}(\phi)\otimes\mathcal{O}_{S}(-C_{0})\ {\rm for\ all}\ \phi\in\mathbb{R}\end{split}\right\},$
(2.7) $\textrm{Stab}(\mathcal{D}_{2})=\left\{(Z_{2},\mathcal{P}_{2})\ \middle|\ \begin{split}&(Z,\mathcal{P})\in\textrm{Stab}(C),Z_{2}=Z\circ F_{2}^{-1},\\ &\mathcal{P}_{2}(\phi)=p^{*}\mathcal{P}(\phi)\ {\rm for\ all}\ \phi\in\mathbb{R}.\end{split}\right\},$
@@
(2.8) $F_{1}:K(D^{b}(C))\cong K(\mathcal{D}_{1}),\quad F_{2}:K(D^{b}(C))\cong K(\mathcal{D}_{2})$
@@ 3. Construction
(3.1) $D^{b}(\Sigma_{e})=\langle p^{*}D^{b}(\mathbb{P}^{1})\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0}),p^{*}D^{b}(\mathbb{P}^{1})\rangle.\\$
@@ 3.1. Gluing construction on $D^{b}(\Sigma_{e})$
(3.2) $\textrm{Hom}_{D^{b}(\Sigma_{e})}(p^{*}E_{1}\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0}),p^{*}E_{2}))\cong\textrm{Hom}_{D^{b}(\mathbb{P}^{1})}(E_{1},E_{2}\otimes(\mathcal{O}\oplus\mathcal{O}(-e))).$
@@
(3.3) $E^{\bullet}:=\cdots\to 0\to\mathcal{O}(m_{j^{\prime}-1})^{\oplus n_{j^{\prime}-1}}\to\mathcal{O}(m_{j^{\prime}})^{\oplus n_{j^{\prime}}}\to 0\to\cdots$
@@
(3.4) $\bigoplus_{i\leq 0}\textrm{Ext}^{i}(\mathcal{O}(k-1)[j_{1}+1],E^{\bullet}[j_{2}])=\bigoplus_{i\leq 0}\textrm{Ext}^{i}(\mathcal{O}(k)[j_{1}],E^{\bullet}[j_{2}])=0.$
@@
(3.5) $\textrm{Ext}^{i+j_{2}-(j_{1}+1)}(\mathcal{O}(k-e-1),\mathcal{F})\quad and\quad\textrm{Ext}^{i+j_{2}-j_{1}}(\mathcal{O}(k-e),\mathcal{F})$
@@
(3.6) $\mathcal{O}(k)^{\oplus i-1-n}\to\mathcal{O}(n)[1]\to\mathcal{O}(k-1)[1]^{\oplus i-n}.$
@@
(3.7) $\mathcal{O}(k)\to\mathcal{O}_{x}\to\mathcal{O}(k-1)[1]$
@@
    (3.9) $\sigma_{k,\zeta_{0},\zeta_{1}}:=(Z_{\zeta_{0},\zeta_{1}},\mathcal{A}(k)[j]),$
@@
    (3.10) $Z_{\zeta_{0},\zeta_{1}}(\mathcal{O}(k-1)[j+1])=\zeta_{0},\quad Z_{\zeta_{0},\zeta_{1}}(\mathcal{O}(k)[j])=\zeta_{1}.$
@@
(3.11) $r_{\lambda_{1}}=\textrm{ch}_{0}(\lambda_{1}(F)),\quad d_{\lambda_{1}}=\textrm{ch}_{1}(\lambda_{1}(F)),\quad r_{\rho_{2}}=\textrm{ch}_{0}(\rho_{2}(F)),\quad d_{\rho_{2}}=\textrm{ch}_{1}(\rho_{2}(F)).$
@@
Next, we show that these four types of pre-stability conditions satisfy the support property. To this end, we first require some preliminary results for the cases $m=1,2,3,4$. In the cases $m=1$ and $3$ (where $\sigma_{2}$ is a quiver stability condition), the following criterion applies.
@@
(3.12) $|Z_{1}(E)|\leq C|Z_{gl,2}(E)|\quad\textrm{and}\quad|Z_{2}(E)|\leq(1+C)|Z_{gl,2}(E)|$
@@
The case $m=4$ is technically more difficult. Unlike the quiver stability case, a lower bound on the phase is not automatically ensured. We postpone the detailed proof to Section 3.2.
@@
For the case $m=4$ follows from the Proposition 3.15. We consider the cases $m=1$ and $3$ (where $\mathcal{A}_{2}$ is the heart of a quiver stability condition). We apply the criterion in Lemma 3.10. Let $E\in\mathcal{A}_{2}$ be a non-zero object. Its central charge is given by $Z_{2}(E)=n_{0}\zeta_{0}+n_{1}\zeta_{1}$, where $\zeta_{0},\zeta_{1}\in\mathbb{H}$, and the dimension vector $(n_{0},n_{1})$ consists of non-negative integers. Since $E\neq 0$, the dimension vector is non-zero, meaning at least one of $n_{0}$ or $n_{1}$ is a positive integer. Observe that the phase of the sum $n_{0}\zeta_{0}+n_{1}\zeta_{1}$ is bounded by the phases of $\zeta_{0}$ and $\zeta_{1}$. We define $\theta:=\min\{\phi(\zeta_{0}),\phi(\zeta_{1})\}$. Since $\zeta_{0},\zeta_{1}\in\mathbb{H}$, the image $Z_{2}(\mathcal{A}_{2}\setminus\{0\})$ is contained in the sector $\mathbb{H}_{\theta}=\{re^{i\pi\phi}\mid\phi\in[\theta,1]\}$. Thus, the condition of Lemma 3.10 is satisfied, and the support property holds.
@@
(3.14) $C=\begin{pmatrix}-1&1\\ 1-k&k\end{pmatrix}$.
@@
(3.15) $C^{-1}=\begin{pmatrix}-k&1\\ 1-k&1\end{pmatrix}$.
@@
(3.16) $\begin{pmatrix}\textrm{rank}\ E\\ \textrm{deg}\ E\end{pmatrix}=C\begin{pmatrix}n_{0}\\ n_{1}\end{pmatrix}=\begin{pmatrix}n_{1}-n_{0}\\ (1-k)n_{0}+kn_{1}\end{pmatrix}$.
@@
(3.17) $\begin{pmatrix}n_{0}\\ n_{1}\end{pmatrix}=\begin{pmatrix}-k&1\\ 1-k&1\end{pmatrix}\begin{pmatrix}r_{*}\\ d_{*}\end{pmatrix}=\begin{pmatrix}d_{*}-kr_{*}\\ d_{*}+(1-k)r_{*}\end{pmatrix}$.
@@
(3.18) $+(-1)^{j_{2}}((d_{\rho_{2}}-kr_{\rho_{2}})\zeta^{\prime}_{0}+(d_{\rho_{2}}+(1-k)r_{\rho_{2}})\zeta^{\prime}_{1}).$
3.2. The support property of $m=4$
In this section, we establish the support property for the glued type $m=4$. Recall that for any object $E\in\mathcal{A}_{gl,4}$, there exist $E_{1},E_{2}\in{\rm Coh}(\mathbb{P}^{1})$ and a canonical short exact sequence
(3.19) $0\to\rho_{2}(E)=p^{*}E_{2}\to E\to p^{*}E_{1}\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0})[1]=\lambda_{1}(E)\to 0$
(cf. (2.3)). Combining the additivity of Chern characters with this sequence yields
$\textrm{ch}(E)$ $=\textrm{ch}(p^{*}E_{2})+\textrm{ch}(p^{*}E_{1}\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0})[1])$
$=(r_{2}-r_{1},r_{1}C_{0}+(d_{2}-d_{1})f,d_{1}+\frac{1}{2}er_{1})$
(3.20) $=:(r,c_{1},\textrm{ch}_{2})$
where $r_{i}:=\textrm{rank}(E_{i})$ and $d_{i}:=\textrm{deg}(E_{i})$ for $i=1,2$. On the other hand, from Proposition 3.7 (4), Proposition 3.11, and Remark 3.5, the stability function $Z_{gl,4}$ can be written as
$Z_{gl,4}$ $=(-1)^{j_{1}}(-d_{\lambda_{1}}+ir_{\lambda_{1}})+(-1)^{j_{2}}(-d_{\rho_{2}}+ir_{\rho_{2}})$
$=d_{\lambda_{1}}-d_{\rho_{2}}+i(r_{\rho_{2}}-r_{\lambda_{1}})$
$=\left(-\textrm{ch}_{2}+\frac{1}{2}ec_{1}.f\right)-\left(\textrm{ch}_{2}+c_{1}.C_{0}+\frac{1}{2}ec_{1}.f\right)+i\left(c_{1}.f+r-(-c_{1}.f)\right)$
(3.21) $=-(2\textrm{ch}_{2}+c_{1}C_{0})+i(r+2c_{1}.f).$
Here, we used the relation $j_{1}=j_{2}+1$ (Remark 3.5) for the second equality. Substituting the Chern characters from (3.20) into (3.21), we obtain the simplified equation:
(3.22) $Z_{gl,4}(E)=-(d_{1}+d_{2})+i(r_{1}+r_{2})=Z_{st}(E_{1})+Z_{st}(E_{2}).$
This is expressed as the sum of two standard stability functions $Z_{st}=-\textrm{deg}+i\textrm{rank}$ on the heart ${\rm Coh}(\mathbb{P}^{1})$. Hence we define
(3.23) $\mu(E):=\frac{d_{1}+d_{2}}{r_{1}+r_{2}}$
and let us define the values
$D:=d_{1}+d_{2}=\textrm{Re}Z_{gl,4}(E),\qquad R:=r_{1}+r_{2}=\textrm{Im}Z_{gl,4}(E).$
From the above, we see that $\mu(\lambda_{1}(E))=\frac{d_{1}}{r_{1}}$ and $\mu(\rho_{2}(E))=\frac{d_{2}}{r_{2}}$ coincide with the slopes $\mu(E_{1})$ and $\mu(E_{2})$ on the base curve. Accordingly, we denote the maximal and minimal slopes by $\mu_{\max}$ and $\mu_{\min}$, respectively.
From equation (3.22), the absolute value of $Z_{gl,4}$ satisfies
(3.24) $|Z_{gl,4}(E)|=\sqrt{(d_{1}+d_{2})^{2}+(r_{1}+r_{2})^{2}}\ \geq\ |d_{1}+d_{2}|,\ r_{1}+r_{2}.$
In order to prove the support property for the case $m=4$, we first require some preliminary results.
Lemma 3.13.
Let $0\neq E\in\mathcal{A}_{gl,4}$ be a $\sigma_{gl,4}$-semistable object and write it as in (3.19). Then:
  1. (1)
    If $R>0$, then both $E_{1}$ and $E_{2}$ are torsion-free, hence vector bundles on $\mathbb{P}^{1}$.
  2. (2)
    If $R=0$, then $E_{1}$ and $E_{2}$ are torsion sheaves and $Z_{gl,4}(E)=-(d_{1}+d_{2})\in\mathbb{R}_{<0}$.
Proof.
(1) Assume $R>0$. If $E_{2}$ has a nonzero torsion subsheaf $T\subset E_{2}$, then $p^{*}T\subset p^{*}E_{2}=\rho_{2}(E)\subset E$ is a subobject in $\mathcal{A}_{gl,4}$. Since $Z_{st}(T)\in\mathbb{R}_{<0}$, we have $p^{*}T\in\mathcal{P}(1)$, hence $\phi(p^{*}T)=1$. On the other hand, $\textrm{Im}Z_{gl,4}(E)=R>0$ implies $\phi^{+}(E)<1$. Thus $\phi(p^{*}T)>\phi^{+}(E)$, contradicting the $\sigma_{gl,4}$-semistability of $E$.
Next, suppose $E_{1}$ has a nonzero torsion subsheaf $T\subset E_{1}$. Then $p^{*}T\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0})[1]$ is a subobject of $\lambda_{1}(E)$ in $\mathcal{A}_{gl,4}$. Since $\rho_{2}(p^{*}T\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0})[1])=0$, this object lifts to a subobject $S\subset E$ in $\mathcal{A}_{gl,4}$. The central charge of $S$ is $Z_{gl,4}(S)=Z_{st}(T)=-\mathrm{deg}(T)\in\mathbb{R}_{<0}$, which implies $\phi(S)=1$. On the other hand, since $\Im Z_{gl,4}(E)=R>0$, we have $\phi(E)<1$. The existence of a subobject $S$ with $\phi(S)>\phi(E)$ contradicts the semistability of $E$. Thus, $E_{1}$ and $E_{2}$ must be torsion-free. On $\mathbb{P}^{1}$, they are vector bundles.
(2) If $R=0$, then $r_{1}=r_{2}=0$, so $E_{1}$ and $E_{2}$ are torsion, and $Z_{gl,4}(E)=-(d_{1}+d_{2})\in\mathbb{R}_{<0}$ follows from the definition. ∎
Lemma 3.14.
Let $0\neq E\in\mathcal{A}_{gl,4}$ be $\sigma_{gl,4}$-semistable and write it as in (3.19). Assume $R=r_{1}+r_{2}>0$, so $E_{1}$ and $E_{2}$ are vector bundles by Lemma 3.13. Let $\mu(E)=\frac{d_{1}+d_{2}}{r_{1}+r_{2}}$ defined in (3.23). Then the following hold.
  1. (1)
    One has the inequality
    (3.25) $\mu_{\max}(E_{2})\ \leq\ \mu(E)\ \leq\ \mu_{\min}(E_{1}).$
  2. (2)
    If $E$ is non-split and $e>0$, then $\textrm{Hom}_{D^{b}(\mathbb{P}^{1})}(E_{1},E_{2})\neq 0$.
  3. (3)
    In either case (split or non-split), $E_{1}$ is slope-semistable and
    (3.26) $d_{1}=\mu(E)\,r_{1}=\frac{r_{1}}{r_{1}+r_{2}}\,(d_{1}+d_{2}).$
    In particular,
    (3.27) $|d_{1}|\leq|d_{1}+d_{2}|,\qquad|d_{2}|\leq 2|d_{1}+d_{2}|.$
Proof.
(1) The existence of the subobject $\rho_{2}(E)\subset E$ and the quotient $E\twoheadrightarrow\lambda_{1}(E)$ forces the phase inequalities. Since $Z_{st}=-\textrm{deg}+i\textrm{rank}$ on $\mathbb{P}^{1}$, these inequalities are equivalent to the HN-inequalities for slope.
(2) By the projection formula and canonical isomorphism $\textbf{R}p_{*}\mathcal{O}_{\Sigma_{e}}(C_{0})\cong\mathcal{O}\oplus\mathcal{O}(-e)$, we have an isomorphism
$\textrm{Ext}^{1}_{\mathcal{A}_{gl,4}}(\lambda_{1}(E),\rho_{2}(E))$ $\cong\textrm{Hom}_{D^{b}(\mathbb{P}^{1})}(E_{1},E_{2}\otimes(\mathcal{O}\oplus\mathcal{O}(-e))$
$\cong\textrm{Hom}_{D^{b}(\mathbb{P}^{1})}(E_{1},E_{2})\oplus\textrm{Hom}_{D^{b}(\mathbb{P}^{1})}(E_{1},E_{2}(-e)),$
where the first isomorphism is given by (3.2). Suppose $\textrm{Hom}(E_{1},E_{2}(-e))\neq 0$. Then by the stability of vector bundles on curves, we have $\mu_{\min}(E_{1})\leq\mu_{\max}(E_{2}(-e))=\mu_{\max}(E_{2})-e$. Combining this with (1), we obtain
$\mu_{\max}(E_{2})\leq\mu(E)\leq\mu_{\min}(E_{1})\leq\mu_{\max}(E_{2})-e.$
This implies $e\leq 0$, which contradicts the assumption $e>0$. Consequently, $\textrm{Hom}(E_{1},E_{2}(-e))=0$. Since $E$ is non-split, the extension class implies a non-zero element in the remaining summand, i.e., $\textrm{Hom}(E_{1},E_{2})\neq 0$.
(3) First assume $E$ is non-split and $e>0$. From (2) we have $\textrm{Hom}(E_{1},E_{2})\neq 0$, which implies $\mu_{\min}(E_{1})\leq\mu_{\max}(E_{2})$. Combining this with the inequality from (1), we obtain a chain of inequalities:
$\mu_{\max}(E_{2})\leq\mu(E)\leq\mu_{\min}(E_{1})\leq\mu_{\max}(E_{2}).$
Consequently, all inequalities must be equalities: $\mu_{\max}(E_{2})=\mu(E)=\mu_{\min}(E_{1})$. Now, suppose for contradiction that $E_{1}$ is not slope-semistable. Then there exists a destabilizing subsheaf $K\subset E_{1}$ with $\mu_{\min}(K)>\mu(E_{1})\geq\mu(E)$. We define a subobject $S:=p^{*}K\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0})[1]\subset\lambda_{1}(E)$. Consider the pullback extension $E^{\prime}\subset E$ defined by the short exact sequence:
$0\to\rho_{2}(E)\to E^{\prime}\to S\to 0.$
The extension class of this sequence lies in $\textrm{Ext}^{1}(S,\rho_{2}(E))$. By the projection formula, we have
$\textrm{Ext}^{1}(S,\rho_{2}(E))\cong\textrm{Hom}(K,E_{2})\oplus\textrm{Hom}(K,E_{2}(-e)).$
Since $\mu_{\min}(K)>\mu(E)\geq\mu_{\max}(E_{2})$, both Hom-groups vanish. Thus, the extension splits, i.e., $E^{\prime}\cong\rho_{2}(E)\oplus S$. This implies that $S$ is a subobject of $E^{\prime}$ and hence of $E$. Since $\phi(S)=\mu(K)>\mu(E)$, this contradicts the semistability of $E$. Hence, $E_{1}$ is slope-semistable of slope $\mu(E)$. The statement (3.26) follows from the slope-semistability of $E_{1}$, which implies (3.27).
If $E$ is split, we have an isomorphism $E\simeq p^{*}E_{2}\oplus(p^{*}E_{1}\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0})[1])$ in $\mathcal{A}_{gl,4}$. Since $E$ is $\sigma_{gl,4}$-semistable, its direct summands must also be $\sigma_{gl,4}$-semistable with the same phase $\phi(E)$. In particular, $\phi(p^{*}E_{1}\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0})[1])=\phi(E)$, which implies $\mu(E_{1})=\mu(E)$. Furthermore, the semistability of $p^{*}E_{1}\otimes\mathcal{O}_{\Sigma_{e}}(-C_{0})[1]$ in $\mathcal{A}_{gl,4}$ implies the slope-semistability of $E_{1}$ on $\mathbb{P}^{1}$, as any destabilizing subsheaf of $E_{1}$ would induce a destabilizing subobject in $\mathcal{A}_{gl,4}$. Thus, $E_{1}$ is slope-semistable with slope $\mu(E)$, and the equalities (3.26) and (3.27) hold. ∎
Now, we fix the norm on $\Lambda_{\mathbb{R}}=K_{num}\otimes\mathbb{R}$ as
$\|\textrm{ch}(G)\|:={\rm max}\left\{|r|,|a|,|b|,|\textrm{ch}_{2}|\right\}$
for $G\in D^{b}(\Sigma_{e})$, where $c_{1}(G)$ is denoted by $aC_{0}+bf$. Using this with the formula (3.20), we obtain
$\|\textrm{ch}(E)\|={\rm max}\left\{|r_{2}-r_{1}|,|r_{1}|,|d_{2}-d_{1}|,\left|d_{1}+\frac{1}{2}er_{1}\right|\right\}$
for all $E\in\mathcal{A}_{gl,4}$.
Proposition 3.15 (Support property for $m=4$).
The glued pre-stability condition $\sigma_{gl,4}=(Z_{gl,4},\mathcal{A}_{gl,4})$ satisfies the support property. Namely, there exists a constant $C>0$ such that for every $\sigma_{gl,4}$-semistable object $0\neq E\in\mathcal{A}_{gl,4}$, we have
$\|\textrm{ch}(E)\|\ \leq\ C\,|Z_{gl,4}(E)|.$
Proof.
Let $E$ be as in (3.19) and set $D=d_{1}+d_{2}$, $R=r_{1}+r_{2}$ defined in (3.23). If $R=0$, then $r_{1}=r_{2}=0$ and $E_{1},E_{2}$ are torsion. In this case $|Z_{gl,4}(E)|=|D|$. Using (3.20) with $r_{1}=r_{2}=0$, we have $r=0$, $a=0$, $b=d_{2}-d_{1}$, and $\textrm{ch}_{2}=d_{1}$. Hence,
$\|\textrm{ch}(E)\|=\max\{|d_{2}-d_{1}|,\ |d_{1}|\}\leq|d_{1}|+|d_{2}|=|D|=|Z_{gl,4}(E)|,$
and hence the inequality holds with constant 11.
Assume now $R>0$. Then $E_{1},E_{2}$ are vector bundles, and Lemma 3.14 holds. In particular, we have $|d_{1}|\leq|D|$ and $|d_{2}|\leq 2|D|$. Using (3.20), the Chern character is given by
$\textrm{ch}(E)=(r,aC_{0}+bf,\textrm{ch}_{2}),\quad r=r_{2}-r_{1},\ a=r_{1},\ b=d_{2}-d_{1},\ \textrm{ch}_{2}=d_{1}+\frac{e}{2}r_{1}.$
We estimate each component in terms of $|Z_{gl,4}(E)|$:
  • $|r|=|r_{2}-r_{1}|\leq R\leq|Z_{gl,4}(E)|$.
  • $|a|=r_{1}\leq R\leq|Z_{gl,4}(E)|$.
  • $|b|=|d_{2}-d_{1}|\leq|d_{1}|+|d_{2}|\leq 3|D|\leq 3|Z_{gl,4}(E)|$.
  • $|\textrm{ch}_{2}|=\left|d_{1}+\frac{e}{2}r_{1}\right|\leq|d_{1}|+\frac{e}{2}|r_{1}|\leq|D|+\frac{e}{2}R\leq\Bigl(1+\frac{e}{2}\Bigr)|Z_{gl,4}(E)|$.
Taking $C:=\max\{3,1+\frac{e}{2}\}$, the inequality holds for all semistable objects. ∎